# A Scalable Bayesian Spatiotemporal Model for Water Level Predictions using a Nearest Neighbor Gaussian Process Approach

### v1

@@ 6 Limitations
This work proposes a model for predicting water level on rivers in Ireland considering a long-term spatiotemporal data. This approach has been demonstrated to be effective, as evidenced by the promising results. However, several limitations should be noted. As previously mentioned, the model incorporates only precipitation as a predictor. The inclusion of additional hydroclimate predictors such as temperature, wind speed, soil moisture and potential evapotranspiration would potentially improve the predictive performance and reliability of the model. Additionally, the misalignment problem between the hydroclimate factors, and the water level was addressed through Inverse Distance Weighting interpolation. Despite the simplicity and efficacy of this methodology, a complex probabilistic interpolation approach could be investigated to potentially enhance the precision of estimations. However, the development of such techniques falls beyond the scope of this study.
In the proposed model, the spatial dependence is captured through a spatial latent field specified as a NNGP, which fundamentally depends on the covariance structure. The introduced structure is straightforward and effective, ensuring computational feasibility without compromising performance. However, it simplifies the complex spatial dependencies inherent in river networks, where flow directionality and confluences may impact the water level measurements along the river. Addressing this limitation could involve specifying a more sophisticated covariance structure that accounts for the distance along the river network and flow directionality. Such covariance structures combined with NNGP are not present in the literature, as far as we are aware, and therefore we view this as a future research direction.
@@ 7 Conclusion
In this paper, we proposed a spatiotemporal NNGP model for predicting water level in a long-term dataset across both space and time. This model not only delivers accurate predictions but also quantifies uncertainty for both monitored stations and unobserved locations. By accounting for spatiotemporal dependencies and incorporating precipitation information, the model offers insights of the impact of precipitation on water levels over time, thereby supporting more informed decision-making in hydrological monitoring, early warning systems, and adaptive resource planning.
@@ 8 Acknowledgments
This publication has emanated from research conducted with the financial support of Science Foundation Ireland under Grant number 18/CRT/6049.

### v2

@@ 6 Limitations
This work proposes a model for predicting water level on rivers in Ireland considering a long-term spatiotemporal data. This approach has been demonstrated to be effective, as evidenced by the promising results. However, several limitations should be noted. As previously mentioned, the model incorporates only precipitation as a predictor. The inclusion of additional hydroclimate predictors such as temperature, wind speed, soil moisture and potential evapotranspiration would potentially improve the predictive performance and reliability of the model. Additionally, the spatial misalignment problem between the observed precipitation, and the water level measurements was addressed using Inverse Distance Weighting interpolation. Despite the simplicity and efficacy of this methodology, more sophisticated probabilistic interpolation could be investigated to potentially enhance the precision of estimations. Nevertheless, the development of such techniques is not the focus of this study, as the misalignment problem introduces additional methodological and computational complexities, and constitutes a research area in its own right (Gotway and Young,, 2002; Finley et al.,, 2014).
In the proposed model, the spatial dependence is captured through a spatial latent field specified as a NNGP for a given covariance function, which fundamentally depends on the Euclidean distance. The introduced structure is straightforward and effective, ensuring computational feasibility without compromising performance. However, this formulation simplifies the complex spatial dependencies inherent in river networks, where flow directionality, varying water volume and confluences may impact the water level measurements along the river. Covariance models that explicitly account for these hydrological features have been developed by Hoef et al., (2006); Cressie et al., (2006); Hoef and Erin, (2010). However, implementing these models is labour-intensive and time-consuming, as it requires extensive processing of stream network data, often involving manual steps. This challenge is further intensified in geographical areas characterised by multiple basins and complex stream connections, such as those found in Ireland.
@@ 7 Conclusion
In this paper, we proposed a spatiotemporal NNGP model for predicting water level in a long-term dataset across both space and time. This model not only delivers accurate predictions but also quantifies uncertainty for both monitored stations and unobserved locations. By accounting for spatiotemporal dependencies and incorporating precipitation information, the model offers insights of the impact of precipitation on water levels over time, thereby supporting more informed decision-making in flood mitigation strategies and resource planning.
Although the main attention in water level prediction in Ireland is predicting for the monitoring location at future time point, since they are strategically placed in areas susceptible to flooding or critical for hydrological control, there is considerable interest in extending predictions to surrounding, unmonitored regions.
@@
As a future research direction, the model could be extended by applying the sparsity-inducing NNGP prior not only to the spatial latent field defined over Euclidean space, but also to the covariance structures for stream networks presented in Hoef and Erin, (2010). The proposed extension would improve the modelling of complex spatial autocorrelation along stream segments and allow Bayesian inference for large hydrological datasets without compromising its predictive performance. To the best of our knowledge, this approach has not yet been explored in the literature.
@@ 8 Acknowledgments
This publication has emanated from research conducted with the financial support of Science Foundation Ireland under Grant number 18/CRT/6049 and the jointly funded by Taighde Éireann - Research Ireland, and GSI under Grant number 20/FFP-P/8610.